Gomi Waste Disposal is evaluating a project that would require the purchase of a piece of equipment for 164,000 dollars today. During year 1, the project is expected to have relevant revenue of 118,000 dollars, relevant costs of 38,000 dollars, and relevant depreciation of 28,000 dollars. Gomi Waste Disposal would need to borrow 164,000 dollars today to pay for the equipment and would need to make an interest payment of 6,000 dollars to the bank in 1 year. Relevant net income for the project in year 1 is expected to be 45,266 dollars. What is the tax rate expected to be in year 1? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

Answer:

Let us calculate Gomi Waste Disposal's Income before tax for Year 1 based on given information:

$118,000 $38,000 $28,000 Revenue Costs Depreciation Total operating expenses Operating Income Interest Income before Tax $66,

We observe Income before tax = $46,000

Given that net income of year 1 = $45,266

Hence:

Tax = 46000 - 45266 = $734

Expected Tax rate in year 1 = 734 / 46000 = .0160 or 1.60%

Expected Tax rate in year 1 = .0160